You are here:
NRAO Public Wiki
TGBT14A_912_75 - continued bank discrepancy tests
Continue bank comparison tests
Ray switches to latest version of Manager
roaches E and H are still swapped.
bank_gain_mode4_kfpa in tp_nocal mode
for some reason which I don't understand, we seem to have much less IF power than before? Set IF target to 7 in balance. Doesn't work
Manually set IF rack atten to 2dB for rcvr Array 18 - 26 J6.
lbw_gain was 163321889, = 2**27.3
Use SetValues to set it to 2**27 = 134217728
both polarizations look good this time. Can use all data.
Scan 1 aborts due to Antenna being in Scan Coordinator
Scans 2 - 17
Repeat with switch mode tp
Run separate 60 sec scans with Cal Off, Cal On
Scan 34 - track with Cal off.
Scan 35 - Cal On
Scan 36 - Cal Off
Scan 37 - Cal On
Scan 38 - Cal Off
Hand over to Ray for snapblock testing at 10:00am.
Galen was putting some absorber over the feed while this was going on, so may have screwed up our tests (We are using B1)....
Get Jason and Ray to swap the Roaches back, but still BOF 123...
Run the Cal on / Cal Off scans with Roaches in normal configuration. Now have absorber over the feed.
config, balance, set values, etc.
Scan 39 - off
Scan 40 - on. Don't see the cal. Maybe system temperature is now way too high?
Try a "tp" scan - 41. GBTIDL can calibrate it, so try the separate scans. Tsys = 370K this actually might be a useful test.
Scan 42 - off
Scan 43,44,45,46 - on, off, on, off
Switch back to normal mode...
scan 47 - running Bank E from Ray's machine
scan 48 - running normally - filler is not running; restart
scan 49. looks fine.
iew wiki text
ore topic actions
Topic revision: r2 - 2015-02-26,
Create New Topic
Copyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki?